# Universal Scaling Laws of Absorbing Phase Transitions in Artificial Deep Neural Networks

### v2

@@ Abstract
We demonstrate that conventional artificial deep neural networks operating near the phase boundary of the signal propagation dynamics—also known as the edge of chaos—exhibit universal scaling laws of absorbing phase transitions in non-equilibrium statistical mechanics. Our numerical results indicate that the multilayer perceptrons and the convolutional neural networks belong to the mean-field and the directed percolation universality classes, respectively. Also, the finite-size scaling is successfully applied, suggesting a potential connection to the depth-width trade-off in deep learning. Furthermore, our analysis of the training dynamics under gradient descent reveals that hyperparameter tuning to the phase boundary is necessary but insufficient for achieving optimal generalization in deep networks. Remarkably, nonuniversal metric factors associated with the scaling laws are shown to play a significant role in concretizing the above observations. These findings highlight the usefulness of the notion of criticality for analyzing the behavior of artificial deep neural networks and offer new insights toward a unified understanding of an essential relationship between criticality and intelligence.
@@ I Introduction
Critical phenomena at second-order phase transitions have long been hypothesized to be the key to the extraordinary computational power of living systems [1, 2, 3]. The idea behind the hypothesis is that information cannot propagate through ordered states of a matter, and it rapidly decays to random noise in the disordered states due to the overly enhanced capability of the medium to convey disturbance [2]. Notably, the notion of phases of matter lies at the core of the hypothesis; this viewpoint focuses on collective aspects of the systems, complementary to more traditional reductionist ones [4, 5]. Despite the experimental and theoretical challenges stemming from the many-body nature of the problem, pursuing the computation at the criticality hypothesis has proven to be a fruitful research direction [3, 6]. In particular, brain dynamics has been intensively discussed [7, 8] in the context of absorbing phase transitions [9, 10] due to the theoretical relation [11] to self-organized criticality [12, 13], not to mention the straightforward correspondence between death and an absorbing state.
@@
Figure 1: Order-to-chaos transition in the multilayer perceptions (1) and the associated universal scaling laws. (a) The phase diagram of the signal propagation for $\tanh$ activation. The solid curve indicates the phase boundary. The dashed line indicates $\sigma_{b}=0.3$, where the maximum Lyapunov exponent $\lambda_{1}$ (see Eq. (3)) and the order parameter $\rho^{(l)}$ (see Eq. (20)) are studied in (b) and (c), respectively. (b) $\lambda_{1}$ as a function of the weight parameter $\sigma_{w}$ for various widths $n=1$ (purple), 2 (green), 9 (light blue), 20 (orange), and 50 (red), estimated from the convergence of the finite counterpart of Eq. (3) at $l=10^{5}$. The vertical line indicates the position of the critical point ($\sigma_{w;c}\sim 1.39558$). The black curve is the result expected in the limit of $n\rightarrow\infty$, namely $\lambda_{\mathcal{C}}/2$ (see Eq. (16)). (c) The main panel shows $\rho^{(l)}$ as a function of $l$ for various $\sigma_{w}$ (from 1.35 (blue; the ordered phase) to 1.45 (magenta; the chaotic phase)), calculated from the mean-field theory (6) and (7). The inset shows the stationary value $\rho^{*}$ of the order parameter and the reciprocal correlation depth $\xi_{\|}^{-1}$ (see Eqs. (13) and (14)) as a function of $\sigma_{w}$. The dashed line and the solid line are guides-to-eye for linear onset with a slope of $\gamma_{\leftrightarrow}$ and $\zeta_{\leftrightarrow}:=\gamma_{\leftrightarrow}/\kappa^{1/\nu_{\|}}$, respectively ($\delta\sigma_{w}:=\sigma_{w}-\sigma_{w;c}$). (d) The order parameter $\rho^{(l)}$ as a function of $l$ for various activation functions is rescaled according to the universal scaling ansatz (17). Different symbols with the same color correspond to different values of $\sigma_{w}$: $+$, $\times$, *, $\square$ in ascending order ($\sigma_{b}$ is fixed to be 0.3). In (c) and (d), the nonuniversal metric factors $\gamma_{\leftrightarrow}$ and $\kappa$ are calculated from Eqs. (22) and (23), respectively, for each case.
@@ II Preliminaries
Here, $\boldsymbol{x}$ is the input with $n_{\mathrm{in}}$ elements, $W^{(l)}$ and $\boldsymbol{b}^{(l)}$ the weight matrix and the bias vector at the $l$-th hidden layer, respectively, $\sigma_{w},\sigma_{b}$ the associated hyperparameters and $n$ the width of the hidden layers. Every element of the weight and the bias is initialized according to the standard normal distribution $\mathcal{N}(0,1)$. This parameterization differs slightly from the one commonly used in practice [31], but the difference is not essential for our study [32]. The activation function $h$ is assumed to be in the $K^{*}=0$ universality class in the sense of Roberts et al. [23], a few examples being $\tanh$, $\mathrm{erf}$, and $\sin$, but not $\mathrm{ReLU}$ (note, however, that similar observations can be made also for ReLU-like activation functions; see Appendix B).
Since the forward dynamics (1) is fully deterministic after initialization, a pair of signals may end up with a collapse: once the two signals $\boldsymbol{z}_{1}^{(l)},\boldsymbol{z}_{2}^{(l)}$ become identical at one hidden layer, they never deviate from each other again at deeper hidden layers. Interpreting the network depth as a temporal degree of freedom, the ordered state $\boldsymbol{z}_{1}^{(l)}=\boldsymbol{z}_{2}^{(l)}$ can be regarded as an absorbing state of the dynamics. Even though the exact order, which requires an accidentally degenerate weight matrix, rarely occurs in practice, one can empirically see cases where the difference between two signals decays exponentially, especially when the network is narrow. However, with sufficient width and a suitable choice of the hyperparameters $(\sigma_{w},\sigma_{b})$, one can observe the opposite: magnification of the difference, even if initially tiny. In this case, the difference no longer converges to a unique $l\rightarrow\infty$ limit but instead fluctuates randomly. Thus, the networks seem to exhibit a phase transition between a unique absorbing (ordered) phase and a fluctuating active (chaotic) phase. As we will see shortly, the transition can be further formalized by studying the linear stability of the exactly ordered state.
@@ III Absorbing property of
By doing so, we can directly see how the notion of the order-to-chaos transition emerges as a many-body effect in the neural networks; see the numerical results in Fig. 1(b). In the case where the hidden layer consists of only a small number $n$ of neurons (say $n\lesssim 10$), the maximum Lyapunov exponent $\lambda_{1}$ as a function of the weight initialization $\sigma_{w}$ is negative in the entire domain, which suggests that the ordered state is always stable against infinitesimal discrepancy. However, $\lambda_{1}$ increases as $n$ becomes larger, and eventually, $\lambda_{1}$ changes its sign at some $\sigma_{w}$ for large $n$, indicating loss of linear stability. Naturally, the position of the onset of the linear instability is very close to that of the critical point predicted from the mean-field theory [17] when $n$ is large and is expected to coincide with the limit of $n\rightarrow\infty$.
@@ IV.1 Mean-field theory of signal propagation
Let us recall some basic results of this theory. One can see that $q^{(l)}$ rapidly converges to a fixed point $q^{*}:=\lim_{l\rightarrow\infty}q^{(l)}$ as the depth $l$ tends to infinity [17, 18], generally without a sign of a phase transition (unless $\sigma_{b}=0$, where $q^{*}$ vanishes at the ordered phase). By substituting $q_{1}^{(l)}=q_{2}^{(l)}=q^{*}$, we obtain an approximate closed-form description for $c^{(l)}$ also known as the iterative $\mathcal{C}$-map, valid for large $l$:
@@
equals to the maximum Lyapunov exponent $\lambda_{1}$ of the ordered state (see Eq. (3)) in the limit of infinite width $n$. It is encouraging to see that the behavior of $\lambda_{1}$ for $n=50$ as a function of $\sigma_{w}$ is already close to the $n\rightarrow\infty$ limit (Fig. 1(b)); this suggests that the infinitely wide neural network may serve as a good starting point for understanding the behavior of the neural networks of practical width.
@@ IV.2 Scaling results for the infinitely wide networks
Figure 2: An intuitive meaning of the nonuniversal metric factor $\kappa$ and its consequence on neural tangent kernel (NTK). (a) The order parameter $\rho^{(l)}$ at a critical point (specifically $(\sigma_{w},\sigma_{b})\sim(1.23367,0.3)$ with erf activation; $\kappa\sim 0.252674$) for various cosine distances $\rho^{(0)}$ (see Eq. (26)) of the inputs, calculated from the mean-field theory (6), (7) and rescaled according to the scaling ansatz (27). The inputs $\boldsymbol{x}_{1},\boldsymbol{x}_{2}$ with $n_{\text{in}}=10$ elements were normalized so that $\|\boldsymbol{x}_{1}\|_{2}=\|\boldsymbol{x}_{2}\|_{2}=1$. The dashed lines are guides-to-eye for the asymptotic behavior (28) of the scaling function $g$. The raw $\rho^{(l)}$ is shown in the inset with a guide to eye for $\rho^{(l)}=(\kappa l)^{-1}$. (b) The NTK $\Theta^{(L)}(\boldsymbol{x}_{1},\boldsymbol{x}_{2})$ (see Eq. (32)) as a function of the network depth $L$, rescaled according to the scaling ansatz (33). The same $(\sigma_{w},\sigma_{b})$ and activation function as (a) are used, and different color corresponds to different $\rho^{(0)}$ of the two inputs. The two horizontal lines are guides-to-eye for the asymptotic behavior (35) as $L\rightarrow\infty$: the upper one for $\boldsymbol{x}_{1}=\boldsymbol{x}_{2}$ ($\rho^{(0)}=0$), the lower one for $\boldsymbol{x}_{1}\neq\boldsymbol{x}_{2}$ ($\rho^{(0)}>0$).
@@
of the initialized networks (here, $y(\boldsymbol{x};\boldsymbol{\theta})$ is the output and $\boldsymbol{\theta}_{0}$ is an initial value of the trainable parameters, namely $\{W^{(l)},\boldsymbol{b}^{(l)}\}_{l=1}^{L+1}$). In the limit of infinite width, the initial value of NTK is deterministic despite the randomness of $\boldsymbol{\theta}_{0}$ themselves, and also NTK stays constant during the training under gradient descent using mean squared error with small learning rate [32]. Consequently, the dynamics of the output $y(\boldsymbol{x};\boldsymbol{\theta})$ under such circumstances can be reduced to a linear ordinary differential equation. In particular, a collection of the residual errors $\Delta\boldsymbol{y}(\boldsymbol{\theta}):=(\Delta y(\boldsymbol{x}_{1};\boldsymbol{\theta}),\cdots,\Delta y(\boldsymbol{x}_{D};\boldsymbol{\theta}))^{T}$ for training inputs $\{\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{D}\}$ is governed by
$\frac{\mathrm{d}\Delta\boldsymbol{y}(\boldsymbol{\theta}(t))}{\mathrm{d}t}=-\frac{2\eta}{D}\Theta_{\mathrm{train}}^{(L)}\Delta\boldsymbol{y}(\boldsymbol{\theta}(t)),$ (30)
@@
$\Theta_{\mathrm{train}}^{(L)}:=\left(\begin{array}[]{ccc}\Theta^{(L)}(\boldsymbol{x}_{1},\boldsymbol{x}_{1})&\cdots&\Theta^{(L)}(\boldsymbol{x}_{1},\boldsymbol{x}_{D})\\ \vdots&\ddots&\vdots\\ \Theta^{(L)}(\boldsymbol{x}_{D},\boldsymbol{x}_{1})&\cdots&\Theta^{(L)}(\boldsymbol{x}_{D},\boldsymbol{x}_{D})\\ \end{array}\right).$ (31)
This makes the initial value of NTK relevant for understanding the training dynamics.
@@
Notice that the term in the product converges to $e^{-{1/\xi_{\|}}}$ in the limit of $l^{\prime}\rightarrow\infty$ unless $\boldsymbol{x}_{1}=\boldsymbol{x}_{2}$, in which case it converges to exponential of the maximum Lyapunov exponent $\lambda_{\mathcal{C}}$ (see Eq. (16)) of the trivial fixed point of the iterative $\mathcal{C}$-map. Taking account of the asymptotic proportionality of the NTK to depth $L$ [19], the form of the scaling ansatz (27) for the order parameter motivates us to consider a universal scaling ansatz for NTK at a critical point:
$\Theta^{(L)}(\boldsymbol{x}_{1},\boldsymbol{x}_{2})\simeq q_{c}^{*}L\tilde{g}(\omega\rho^{(0)}\kappa L),$ (33)
where $\tilde{g}$ is a suitable scaling function and $\rho^{(0)}$ is the cosine distance of the two inputs $\boldsymbol{x}_{1},\boldsymbol{x}_{2}$. Reasonable scaling collapse (except when $L$ is small, just as expected) shown in Fig. 2(b) validates the ansatz. We can also see that the scaling function $\tilde{g}$ satisfies the following asymptotic behavior:
$\tilde{g}(x)\simeq\begin{cases}1&x\ll 1;\\ 1/3&x\gg 1.\end{cases}$ (34)
The scaling form (33), together with the asymptotics (34), suggests that $\kappa L$ is a crucial factor for the properties of NTK at a critical point. If $\kappa L$ is too small, the resulting $\Theta_{\mathrm{train}}^{(L)}$ becomes nearly rank-1, which implies slow training of the network in general. Conversely, if $\kappa L$ is too large, NTK asymptotically behaves like an indicator function [21]
@@
and therefore the dynamics of an output $y(\boldsymbol{x})$ for an input $\boldsymbol{x}$ outside the training dataset, namely,
$\frac{\mathrm{d}y(\boldsymbol{x};\boldsymbol{\theta}(t))}{\mathrm{d}t}=-\frac{2\eta}{D}\Theta_{\text{test}}^{(L)}(\boldsymbol{x})\Delta\boldsymbol{y}(\boldsymbol{\theta}(t))$ (36)
@@
$\Theta_{\text{test}}^{(L)}(\boldsymbol{x}):=(\Theta^{(L)}(\boldsymbol{x},\boldsymbol{x}_{1}),\cdots,\Theta^{(L)}(\boldsymbol{x},\boldsymbol{x}_{D})),$ (37)

### v3

@@ Abstract
We demonstrate that conventional artificial deep neural networks operating near the phase boundary of the signal propagation dynamics—also known as the edge of chaos—exhibit universal scaling laws of absorbing phase transitions in non-equilibrium statistical mechanics. We exploit the fully deterministic nature of the propagation dynamics to elucidate an analogy between a signal collapse in the neural networks and an absorbing state (a state that the system can enter but cannot escape from). Our numerical results indicate that the multilayer perceptrons and the convolutional neural networks belong to the mean-field and the directed percolation universality classes, respectively. Also, the finite-size scaling is successfully applied, suggesting a potential connection to the depth-width trade-off in deep learning. Furthermore, our analysis of the training dynamics under the gradient descent reveals that hyperparameter tuning to the phase boundary is necessary but insufficient for achieving optimal generalization in deep networks. Remarkably, nonuniversal metric factors associated with the scaling laws are shown to play a significant role in concretizing the above observations. These findings highlight the usefulness of the notion of criticality for analyzing the behavior of artificial deep neural networks and offer new insights toward a unified understanding of the essential relationship between criticality and intelligence.
@@ I Introduction
Critical phenomena at second-order phase transitions have long been hypothesized to be the key to the extraordinary computational power of living systems [1, 2, 3]. The idea behind the hypothesis is that information cannot propagate through ordered states of matter, and it rapidly decays to random noise in the disordered states due to the overly enhanced capability of the medium to convey disturbance [2]. Notably, the notion of phases of matter lies at the core of the hypothesis; this viewpoint focuses on collective aspects of the systems, complementary to more traditional reductionist ones [4, 5]. Despite the experimental and theoretical challenges stemming from the many-body nature of the problem, pursuing the computation at the criticality hypothesis has proven to be a fruitful research direction [3, 6]. In particular, brain dynamics has been intensively discussed [7, 8] in the context of absorbing phase transitions [9, 10] due to the theoretical relation [11] to self-organized criticality [12, 13], not to mention the straightforward correspondence between death and an absorbing state.
@@
Figure 1: Order-to-chaos transition in the multilayer perceptions (1) and the associated universal scaling laws. (a) The phase diagram of the signal propagation for $\tanh$ activation. The solid curve indicates the phase boundary. The dashed line indicates $\sigma_{b}=0.3$, where the maximum Lyapunov exponent $\lambda_{1}$ (see Eq. (3)) and the order parameter $\rho^{(l)}$ (see Eq. (20)) are studied in (b) and (c), respectively. (b) $\lambda_{1}$ as a function of the weight parameter $\sigma_{w}$ for various widths $n$, estimated from the convergence of the finite counterpart of Eq. (3) at $l=10^{5}$. The vertical line indicates the position of the critical point ($\sigma_{w;c}\sim 1.39558$). The black curve is the result expected in the limit of $n\rightarrow\infty$, namely $\lambda_{\mathcal{C}}/2$ (see Eq. (16)). (c) The main panel shows $\rho^{(l)}$ as a function of $l$ for various $\sigma_{w}$ (from 1.35 (blue; the ordered phase) to 1.45 (magenta; the chaotic phase)), calculated from the mean-field theory (6) and (7). The inset shows the stationary value $\rho^{*}$ of the order parameter and the reciprocal correlation depth $\xi_{\|}^{-1}$ (see Eqs. (13) and (14)) as a function of $\sigma_{w}$. The dashed line and the solid line are guides-to-eye for linear onset with a slope of $\gamma_{\leftrightarrow}$ and $\zeta_{\leftrightarrow}:=\gamma_{\leftrightarrow}/\kappa^{1/\nu_{\|}}$, respectively ($\delta\sigma_{w}:=\sigma_{w}-\sigma_{w;c}$). (d) The order parameter $\rho^{(l)}$ as a function of $l$ for various activation functions is rescaled according to the universal scaling ansatz (17). Different symbols with the same color correspond to different values of $\sigma_{w}$: $+$, $\times$, *, $\square$ in ascending order ($\sigma_{b}$ is fixed to be 0.3). In (c) and (d), the nonuniversal metric factors $\gamma_{\leftrightarrow}$ and $\kappa$ are calculated from Eqs. (22) and (23), respectively, for each case.
@@ II Preliminaries
Here, $\boldsymbol{x}$ is the input with $n_{\mathrm{in}}$ elements, $W^{(l)}$ and $\boldsymbol{b}^{(l)}$ the weight matrix and the bias vector at the $l$-th hidden layer, respectively, $\sigma_{w},\sigma_{b}$ the associated hyperparameters and $n$ the width of the hidden layers; the meanings of the symbols associated with the neural networks are also summarized in Table 1 for convenience. Every element of the weight and the bias is initialized according to the standard normal distribution $\mathcal{N}(0,1)$. This parameterization differs slightly from the one commonly used in practice [31], but the difference is not essential for our study [32]. The activation function $h$ is assumed to be in the $K^{*}=0$ universality class in the sense of Roberts et al. [23], a few examples being $\tanh$, $\mathrm{erf}$, and $\sin$, but not $\mathrm{ReLU}$ (note, however, that similar observations can be made also for ReLU-like activation functions; see Appendix B).
Since the forward dynamics (1) is fully deterministic after initialization, a pair of signals may end up with a collapse: once the two signals $\boldsymbol{z}_{1}^{(l)},\boldsymbol{z}_{2}^{(l)}$ become identical at one hidden layer, they never deviate from each other again at deeper hidden layers. Interpreting the network depth as a temporal degree of freedom, the ordered state $\boldsymbol{z}_{1}^{(l)}=\boldsymbol{z}_{2}^{(l)}$ can be regarded as an absorbing state of the dynamics. Even though the exact order rarely occurs in practice because it requires an accidentally degenerate weight matrix, the difference between two signals can decay exponentially in some cases, especially when the network is narrow. Meanwhile, with sufficient width and a suitable choice of the hyperparameters $(\sigma_{w},\sigma_{b})$, one can observe the opposite: magnification of the difference, even if initially tiny. In this case, the difference no longer converges to a unique $l\rightarrow\infty$ limit but instead fluctuates randomly. Thus, the networks seem to exhibit a phase transition between a unique absorbing (ordered) phase and a fluctuating active (chaotic) phase. As we will see shortly, the transition can be further formalized by studying the linear stability of the exactly ordered state.
Table 1: Glossary of the symbols associated with the neural networks. If accompanied by the superscript “$(l)$,” the symbol indicates the corresponding quantity at the $l$-th hidden layer.
Symbol Meaning
$\boldsymbol{x}$ Input vectors.
$W$ Weight matrices.
$\boldsymbol{b}$ Bias vectors.
$\sigma_{w},\sigma_{b}$ Hyperparameters associated with the network,
corresponding to standard deviation of the weight/bias.
$n_{\text{in}}$ The number of the elements of the input.
$n$ Width of the network.
$L$ Depth of the network.
$\boldsymbol{z}$ Preactivation: the signal vector just before
applying the activation function.
$h$ Activation function.
$q$ Preactivation variance in the mean-field theory.
$C$ Preactivation covariance in the mean-field theory.
@@ III Absorbing property of
By doing so, we can directly see how the notion of the order-to-chaos transition emerges as a many-body effect in the neural networks. For instance, $\lambda_{1}$ as a function of $\sigma_{w}$ for $\tanh$ activation function is shown in Fig. 1(b). In this case, $\lambda_{1}$ is negative in the entire domain for small width $n$, which suggests that the ordered state is always stable against infinitesimal discrepancy. However, $\lambda_{1}$ increases as $n$ becomes larger, and eventually, $\lambda_{1}$ changes its sign at some $\sigma_{w}$ for large $n$, indicating loss of linear stability. Naturally, the position of the onset of the linear instability is very close to that of the critical point predicted from the mean-field theory [17] when $n$ is large and is expected to coincide with the limit of $n\rightarrow\infty$. One can also empirically see that the loss of the stability of the ordered state at some $\sigma_{w}$ for sufficiently large $n$ is a robust feature of the neural networks under the current consideration, although how $\lambda_{1}$ as a function of $\sigma_{w}$ behaves for small $n$ is somewhat more sensitive to the choice of $h$.
@@ IV.1 Mean-field theory of signal propagation
Let us recall some basic results of this theory. One can see that $q^{(l)}$ rapidly converges to a fixed point $q^{*}:=\lim_{l\rightarrow\infty}q^{(l)}$ as the depth $l$ tends to infinity [17, 18], generally without a sign of a phase transition (unless $\sigma_{b}=0$, where $q^{*}$ vanishes at the ordered phase). By substituting $q_{1}^{(l)}=q_{2}^{(l)}=q^{*}$ to Eqs. (7), (8) and observing Eq. (10), we obtain an approximate closed-form description for $c^{(l)}$ also known as the iterative $\mathcal{C}$-map, valid for large $l$:
@@
equals to the maximum Lyapunov exponent $\lambda_{1}$ of the ordered state (see Eq. (3)) in the limit of infinite width $n$. Encouragingly, the behavior of $\lambda_{1}$ for $n=50$ as a function of $\sigma_{w}$ is already close to the $n\rightarrow\infty$ limit (Fig. 1(b)); this suggests that the infinitely wide neural network may serve as a good starting point for understanding the behavior of the neural networks of practical width.
@@ IV.2 Scaling results for the infinitely wide networks
Figure 2: Critical dynamic scaling of the order parameter $\rho^{(l)}$ and its consequence on neural tangent kernel (NTK). (a) The order parameter $\rho^{(l)}$ at a critical point (specifically $(\sigma_{w},\sigma_{b})\sim(1.23367,0.3)$ with erf activation; $\kappa\sim 0.252674$) for various cosine distances $\rho^{(0)}$ (ranging from $10^{-4}$ (orange) to $10^{-1}$ (purple); see Eq. (26)), calculated from the mean-field theory (6), (7) and rescaled according to the scaling ansatz (27). The inputs $\boldsymbol{x}_{1},\boldsymbol{x}_{2}$ with $n_{\text{in}}=10$ elements were normalized so that $\|\boldsymbol{x}_{1}\|_{2}=\|\boldsymbol{x}_{2}\|_{2}=1$. The dashed lines are guides-to-eye for the asymptotic behavior (28) of the scaling function $g$. The raw $\rho^{(l)}$ is shown in the inset with a guide to eye for $\rho^{(l)}=(\kappa l)^{-1}$. (b) The NTK $\Theta^{(L)}(\boldsymbol{x}_{1},\boldsymbol{x}_{2})$ (see Eq. (32)) as a function of the network depth $L$, rescaled according to the scaling ansatz (35). The same $(\sigma_{w},\sigma_{b},\rho^{(0)})$ and activation function as (a) were used. The two horizontal lines are guides-to-eye for the asymptotic behavior (38) as $L\rightarrow\infty$: the upper one for $\boldsymbol{x}_{1}=\boldsymbol{x}_{2}$ ($\rho^{(0)}=0$), the lower one for $\boldsymbol{x}_{1}\neq\boldsymbol{x}_{2}$ ($\rho^{(0)}>0$).
@@
of the initialized networks (here, $y(\boldsymbol{x};\boldsymbol{\theta})$ is the output and $\boldsymbol{\theta}_{0}$ is an initial value of the trainable parameters, namely $\{W^{(l)},\boldsymbol{b}^{(l)}\}_{l=1}^{L+1}$). In the limit of infinite width, the initial value of NTK is deterministic despite the randomness of $\boldsymbol{\theta}_{0}$ themselves, and also NTK stays constant during the training under the gradient descent using mean squared error with small learning rate [32]. Consequently, the dynamics of the output $y(\boldsymbol{x};\boldsymbol{\theta})$ under such circumstances can be reduced to a linear ordinary differential equation. In particular, a collection of the residual errors $\Delta\boldsymbol{y}(\boldsymbol{\theta}):=(\Delta y(\boldsymbol{x}_{1};\boldsymbol{\theta}),\cdots,\Delta y(\boldsymbol{x}_{N};\boldsymbol{\theta}))^{T}$ for training inputs $\{\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{N}\}$ is governed by
$\frac{\mathrm{d}\Delta\boldsymbol{y}(\boldsymbol{\theta}(t))}{\mathrm{d}t}=-\frac{2\eta}{N}\Theta_{\mathrm{train}}^{(L)}\Delta\boldsymbol{y}(\boldsymbol{\theta}(t)),$ (30)
@@
$\Theta_{\mathrm{train}}^{(L)}:=\left(\begin{array}[]{ccc}\Theta^{(L)}(\boldsymbol{x}_{1},\boldsymbol{x}_{1})&\cdots&\Theta^{(L)}(\boldsymbol{x}_{1},\boldsymbol{x}_{N})\\ \vdots&\ddots&\vdots\\ \Theta^{(L)}(\boldsymbol{x}_{N},\boldsymbol{x}_{1})&\cdots&\Theta^{(L)}(\boldsymbol{x}_{N},\boldsymbol{x}_{N})\\ \end{array}\right).$ (31)
NTK also plays a significant role in characterizing generalizability of the infinitely-wide neural networks trained under the stochastic gradient descent [41]. Thus, the initial value of NTK is relevant for understanding the training dynamics and performance of the neural networks.
@@
The both terms in the RHS of Eq. (32) at a critical point is directly related to the order parameter $\rho^{(l)}$ (to find Eq. (34), perform a similar infinitesimal expansion to the one demonstrated in Appendix A)
$C^{(l)}\simeq q_{c}^{*}(1-\rho^{(l)}),$ (33)
$\sigma_{w;c}^{2}\int\mathcal{D}z_{1}\int\mathcal{D}z_{2}\,h^{\prime}(u_{1}^{(l)})h^{\prime}(u_{2}^{(l)})\simeq 1-2\kappa\rho^{(l)}.$ (34)
As such, we naturally expect a universal scaling ansatz akin to Eq. (27) for the NTK at a critical point:
$\Theta^{(L)}(\boldsymbol{x}_{1},\boldsymbol{x}_{2})\simeq q_{c}^{*}L\tilde{g}(\omega\rho^{(0)}\kappa L)$ (35)
with a suitable scaling function $\tilde{g}$. Here, the prefactor $q_{c}^{*}L$ reflects the asymptotic proportionality to $L$ established in the literature [19, 21]; the limits of the scaling function $\tilde{g}$ can also be deduced from the prior works
$\tilde{g}(x)\rightarrow\begin{cases}1&\text{as}\quad x\rightarrow 0;\\ 1/3&\text{as}\quad x\rightarrow\infty.\end{cases}$ (36)
Alternatively, one may heuristically understand these limits by substituting $\rho^{(l)}=0$ to Eqs. (33), (34) (for $x\rightarrow 0$) or by plugging the asymptote $\rho^{(l)}\simeq(\kappa l)^{-1}$ into Eq. (34) and then resorting to an arithmetic formula (for $x\rightarrow\infty$)
$\lim_{L\rightarrow\infty}\left[\frac{1}{L}\sum_{l=1}^{L}\prod_{l^{\prime}=l}^{L}\left(1-\frac{2}{l^{\prime}}\right)\right]=\frac{1}{3}.$ (37)
Either way, reasonable scaling collapse (except when $L$ is small, just as expected) shown in Fig. 2(b) empirically validates the scaling ansatz (35).
The ansatz (35), together with the asymptotics (36), suggests that $\kappa L$ is a crucial factor for the properties of NTK at a critical point. If $\kappa L$ is too small, the resulting $\Theta_{\mathrm{train}}^{(L)}$ becomes nearly rank-1, which implies slow training of the network in general. Conversely, if $\kappa L$ is too large, NTK asymptotically behaves like an indicator function
@@
which seriously deteoriates the network. For instance, the dynamics of an output $y(\boldsymbol{x})$ for an input $\boldsymbol{x}$ outside the training dataset under the gradient descent, namely,
$\frac{\mathrm{d}y(\boldsymbol{x};\boldsymbol{\theta}(t))}{\mathrm{d}t}=-\frac{2\eta}{N}\Theta_{\text{test}}^{(L)}(\boldsymbol{x})\Delta\boldsymbol{y}(\boldsymbol{\theta}(t))$ (39)
@@
$\Theta_{\text{test}}^{(L)}(\boldsymbol{x}):=(\Theta^{(L)}(\boldsymbol{x},\boldsymbol{x}_{1}),\cdots,\Theta^{(L)}(\boldsymbol{x},\boldsymbol{x}_{N})),$ (40)